OpenStudy (anonymous):

Please explain me this: 1) z²-3z-4 = (z-4)(z+1) 2) 2z²+3z-2 = (2z-1)(z+2)

OpenStudy (anonymous):

1) \[z ^{2} - 3z -4 = z ^{2} -4z+z-4 = z(z-4)+1(z-4) =(z-4)(z+1)\]

OpenStudy (anonymous):